OpenStudy (anonymous):

16! over4! times 12!=

OpenStudy (kainui):

Remember that factorial just means every number counting up to that one multiplied by it. So... 4!=1*2*3*4 Expand it out and see for yourself.

OpenStudy (anonymous):

do u want to know the answer? or how to do it?

OpenStudy (anonymous):

4!=24

OpenStudy (anonymous):

both

OpenStudy (kainui):

Just apply what you just learned, you're right.

OpenStudy (anonymous):

well, 16! is quite a large number

OpenStudy (kainui):

It's not that large to write out on paper.

OpenStudy (anonymous):

Is it (16!/4!)*12! or (16!)/(4!*12)

OpenStudy (kainui):

There's only one way to learn math in my book, and that's by doing. Expanding out the factorial will give firsthand experience with what's going on and he won't have to keep asking us questions. =D

OpenStudy (anonymous):

2.092278989x10^13 is large

OpenStudy (anonymous):

(16!)/(4!*12)

OpenStudy (anonymous):

expand 16! upto 4! thn cncl out 4! over 4! and mutaltiply remaining numbers and thb find 12! and do simple calculation

OpenStudy (anonymous):

A. .0238 B. 1,820 C. 2,380 those are the choices

OpenStudy (anonymous):

since x! = 1*2*3*4..*x: 16! = 12! *13*14*15*16 (12! *13*14*15*16) / (4!*12!) = (13*14*15*16)/(4!) = (43680/(4!) =43680/(1*2*3*4) = 43680/24 = 1820

OpenStudy (kainui):

@uj1024 If you expand it out without simplifying you will see that most of the terms cancel out. It's not that crazy...

OpenStudy (anonymous):

Since since x! = 1*2*3*4..*x: 4! = 1*2*3*4 4! = 2!*3*4 4! = 3!*4 x! is just a way to represent multiplying 1 to x all together (whole numbers only).